OpenStudy (anonymous):

Find the equation of a line in standard form given that it is perpendicular to the line and passes through the point . Graph the new line on an x-y plane given a restricted domain of . Step by Step pls

OpenStudy (anonymous):

What is the point?

OpenStudy (anonymous):

line = 2x-5y=9, point (-1,-1). I don't know what to do after -(1/2)x + By = C

OpenStudy (anonymous):

Alright so rearrange the first line you gave me into y = mx + b format

OpenStudy (anonymous):

domain of -2<x =< 1 =< i mean by less or equal to

OpenStudy (anonymous):

y = -1/2x -1

OpenStudy (anonymous):